Demonstrate your understanding of the following through reference to the relevant authors and the findings from their research.

• Newton’s Laws of Motion.

• Ideal gas laws.

• You must apply both of these to a range of physical situations by giving appropriate examples.


1
Expert's answer
2021-05-05T07:13:41-0400

Answers:-


1.)

Newton’s laws of motion


First law -, if a body is in rest or moving at a constant speed in a straight line than, it will remain at rest or may keep moving in a straight line at constant velocity unless or untill it is acted upon by a external force.

eg. A person sitting in bus



 F = ma

eg. simple example is , we pull or push lighter things easily compared to heavier things.


Third law - to an action there is equal and opposite reaction.

eg. the recoil force by gun is example of third law.


2.) Ideal gas laws

(basic law is pv = nRT , p=pressure , v=volume , n =amount of substanc , R e=ideal gas constant , T=temperature)

Boyle’s Law it is

"P \propto \dfrac{1}{V}"


eg.

"P_1V_1=P_2V_2"


Charles's Law 

"V\propto \; T"


eg.

"\dfrac{V_1}{T_1}=\dfrac{V_2}{T_2}"


Avogadro's Law


"V \propto \; n"


eg.

"\dfrac{V_1}{n_1}=\dfrac{V_2}{n_2}"


Amontons's Law


"P \propto \; T"


eg.

"\dfrac{P_1}{T_1}=\dfrac{P_2}{T_2}"